opnicate Choice of will learn to the electiokyticsolution. PRE-LAB 1. Calculate the theoretical yield of copper that could
Answer #1

souhon Given that I cu 2+ + 2 ē - cu Current C) = 0.50 A = 0.56c/s = 600 sec Hime (t) = 100 min x 60s mio charge (a) = curren

consider the half reachion. 1. The half reach on of an electrolytic cell is - 02+uh tué - 2H₂O e = 1.23 The given elecholytic

teen - Eren - o: osaa log CF 14 - cut2]. Ecen = -0.55- 0.0592 - log [ft] [cuta). fcell to.55 -Eclogs H+7 + Ve log[cutz) o-osa